# SH2: Self-Highlighted Hesitation Helps You Decode More Truthfully

### v3

@@ Abstract
Large language models (LLMs) demonstrate great performance in text generation. However, LLMs are still suffering from hallucinations. In this work, we propose an inference-time method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 is based on a simple fact rooted in information theory that for an LLM, the tokens predicted with lower probabilities are prone to be more informative than others. Our analysis shows that the tokens assigned with lower probabilities by an LLM are more likely to be closely related to factual information, such as nouns, proper nouns, and adjectives. Therefore, we propose to “highlight” the factual information by selecting the tokens with the lowest probabilities and concatenating them to the original context, thus forcing the model to repeatedly read and hesitate on these tokens before generation. During decoding, we also adopt contrastive decoding to emphasize the difference in the output probabilities brought by the hesitation. Experimental results demonstrate that our SH2, requiring no additional data or models, can effectively help LLMs elicit factual knowledge and distinguish hallucinated contexts. Significant and consistent improvements are achieved by SH2 for LLaMA-7b, LLaMA2-7b and Mistral-7b on multiple hallucination tasks.  [^1] We release our code at https://github.com/0-KaiKai-0/SH2.
@@
Jushi Kai  Hai Hu  Zhouhan Linthanks:   Corresponding author Shanghai Jiao Tong University {json.kai, hu.hai}@sjtu.edu.cnlin.zhouhan@gmail.com
@@ 1 Introduction
Depending on massive training corpora, large language models (LLMs) have made tremendous progress in natural language understanding and text generation (Touvron et al., 2023a, Jiang et al., 2023, OpenAI, 2023). However, during reasoning and generation, LLMs could suffer from hallucinations and generate non-factual answers (Zhang et al., 2023b).
Existing approaches address this problem mainly by targeting the generalization error during inference. It has been found that a problematic alignment process could mislead LLMs into hallucination (Lin et al., 2022, Zhang et al., 2023a). In the domain not included in the training, LLMs assign similar probabilities to correct and wrong choices since they do not have enough relevant knowledge to distinguish them.
To fill the gap of knowledge, retrieval augmentation methods Peng et al. (2023); Gao et al. (2023); Gou et al. (2023) leverage external knowledge bases and tools to correct the output of LLMs. Although they provide additional information for LLMs, it still seems impossible to guarantee zero error in the external knowledge. Other researchers propose decoding reformulation methods Li et al. (2023b); O’Brien and Lewis (2023); Chuang et al. (2023) to address hallucinations. However, most of them rely on external data with human labor or powerful models to rescale probability distribution during decoding.
Recent works Wei et al. (2022); Goyal et al. (2023) have demonstrated that a few more computation steps for LLMs can make a difference. Wei et al. (2022) propose chain-of-thought prompting to elicit intermediate reasoning steps of LLMs. These intermediate steps can lead to better answers. However, elaborate prompts need to be prepared in advance. Goyal et al. (2023) append learnable pause tokens to the input prefix and train models from scratch. These delays introduced by the pause tokens provide the model with more computation steps to generate better answers.
In this paper, we seek to leverage the LLM itself to highlight the informative tokens and digest them as input during the hesitation steps to elicit truthful knowledge inside LLMs. We propose a simple yet effective method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 introduces hesitations to give LLMs more time to understand contexts and answer questions. For LLMs, the tokens assigned with lower probabilities are harder to predict, while more likely to be informative. LLMs can select these key tokens by themselves from the input and hesitate on these highlighted tokens. We calculate the difference brought by highlighted tokens through contrastive decoding (Li et al., 2023c) and integrate it into the output probability. Experiments on multiple tasks demonstrate that such a difference could elicit knowledge about the facts inside the model and significantly mitigate hallucinations. Unlike other methods, our method does not leverage any other external tools or data. Additionally, it can be directly deployed during inference with no more training.
@@ 2 Related Work
Since training approaches have high requirements for computational resources, many inference-time approaches are proposed to circumvent LLM hallucinations. These methods can also be divided into retrieval augmentation and decoding reformulation.
@@
Retrieval augmentation approaches resort to external knowledge bases and tools to help correct hallucinations. Additional information is retrieved to provide relevant knowledge for LLMs and support their generation. Peng et al. (2023) and Gao et al. (2023) leverage search engines to attribute and refine the output of language models. Furthermore, Gou et al. (2023) enable multiple tools to correct responses of LLMs autonomously during the interaction with external tools.
@@
Decoding reformulation approaches work on reformulating the probability distribution of the outputs. Li et al. (2023b) introduce Inference-Time Intervention (ITI) to locate truthful directions of TruthfulQA Lin et al. (2022) and shift model activations toward truthfulness during inference. Despite no additional training data, they still need to train a domain-specific classifier for each attention head with the data of TruthfulQA. While Li et al. (2023c) propose Contrastive Decoding (CD) to capture the likelihood difference between large and small models. They suggest that this difference signals which texts should be preferred when generating. And O’Brien and Lewis (2023) leverage contrastive decoding to improve reasoning quality for LLMs. However, the original contrastive decoding must utilize a strong LLM as the expert model. Chuang et al. (2023) propose DoLa to localize the premature layer in LLaMA Touvron et al. (2023a), which contains the least knowledge about the input text. They use the last layer as the expert model and the premature layer as the amateur model, and contrast the prediction probability between them. Shi et al. (2023) leverage contrastive decoding to amplify the output difference between with and without contexts for summarization tasks.
@@
The idea of using extra decoding steps while predicting harder tokens can be dated back to Adaptive Computation Time proposed by Graves (2017). Shin et al. (2020) introduce AUTOPROMPT to combine the original inputs with trigger tokens to elicit knowledge from pretrained models. Wiegreffe et al. (2022) investigate that additional steps to generate rationales could lead to a more faithful model. More recently, Goyal et al. (2023) have also demonstrated that for LLMs, inserting extra computation steps to allow the model to pause before generation can enhance the performance on question answering and reasoning tasks. Nevertheless, their method only works when the model is both pre-trained and finetuned with pauses.
@@
Figure 1: The pipeline to construct and leverage our Self-Highlighted Hesitation.
@@ 3 Self-Highlighted Hesitation
The illustration of our Self-Highlighted Hesitation is shown in Figure 1. For the original inference procedure, the instruction, document, and summary are directly fed into the LLM to generate the judgment. The LLM can be easily confused by the hallucinated context.
For our SH2, we underline the key tokens of the input document by the prediction likelihood. These tokens are listed as a hesitation and appended to the document. The prediction probability is scaled with the difference of confidence between the hesitation-prompted input and the original input. Our method can effectively help LLMs identify the hallucinated context. We will elaborate our method in the following subsections.
@@ 3.2 Relation between Key Tokens and Factual Information
where $POS(\cdot)$ is the function to derive POS tags and #$\#\{POS(X_{i})=z_{k}\}$ is to count how many words in $X_{i}$ has the POS tag of $z_{k}$.
For a dataset with $m$ documents, the recall of POS tag $z_{k}$ can be calculated by:
@@
where $T(X_{i},\eta)$ is the set of words that are among the lowest $\eta$ portion of probability predicted by the language model in $X_{i}$. The $\eta$ in the denominator is to normalize the scale of $\Delta_{\eta}(z_{k})$
$\Delta_{\eta}(z_{k})$ measures how hard the words are with the POS tag $z_{k}$ for LLMs to predict with respect to the proportion $\eta$. $\Delta_{\eta}(z_{k})$ is larger than 1 means that $z_{k}$ is concentrated in the hardest part of documents. Specifically, $\Delta$ will be 1 for all the tags when $\eta$ is set to be 100%. Figure 2 demonstrates the normalized top-$\eta$ recall of the most frequent POS tags for the summarization track of HaluEval Li et al. (2023a).
It can be observed from Figure 2 that content words such as adjectives (JJ), nouns (NN), proper nouns (NNP), adverbs (RB) and conjugated verbs (VBD, VBG, VBP) are more frequent in the hardest words we extract, as the light color in the heatmap indicates. On the other hand, function words such as conjunctions (CC), determiners (DT), and prepositions (IN) are less frequent, as indicated by the darker colors. [^2] The base form of the verbs (VB) seems to be an exception. It is probably because VBs have less information load than conjugated verb forms, as the latter also encode tense, aspect, and other information about the verb. This suggests that the normalized top-$\eta$ recall is positively correlated with factual informativeness. Therefore, picking out key tokens that carry more content, rather than those that serve as grammatical functions, could intuitively help the model to focus on factual information. Our method can thus be understood as a way to “highlight” the content words to LLMs, potentially making them less likely to hallucinate, the effect of which will be evaluated in our experiments. More cases will be given in Appendix D.2.
@@ 3.3 Construction of Hesitations
To this end, we attempt to solve hallucinations by highlighting the key tokens that are hard to predict. For the text sequence $X=(x_{0},x_{1},\dots,x_{n})$, we can obtain the corresponding probability sequence by the language model:
@@
We use the following strategy to select key tokens according to their probability $\hat{P}(X)$ and construct hesitations:
First, construct a candidate key-token set by selecting $\eta\cdot n$ tokens with the lowest prediction probabilities of $\hat{P}(X)$, where $\eta\in(0,1)$ is the sampling proportion we preset for $X$. Then, derive the target key-token set $T(X)$ by randomly drop out $\lambda\cdot\eta\cdot n$ tokens from the candidate set, where $\lambda\in[0,1]$ is the drop-out rate.
For long documents, to avoid $T(X)$ got dominated by one or two POS tags, we introduce the drop-out rate $\lambda$ to randomly select key tokens among a larger pool of candidates, whose size is determined by $\eta$. As a result, $(1-\lambda)\cdot\eta\cdot n$ tokens are retained in $T(X)$ to fill as input during the hesitation steps.
Finally, construct the Hesitation text $H(X)$ with the target key-token set $T(X)$ for the input context: $H(X)=``Pondering:<T(X)>."$, where ``$Pondering:"$ is the prefix. We keep the tokens’ order of $T(X)$ in which they appear in the original text $X$.
@@ 3.4 Contrastive Decoding on Hesitations
Using hesitations as prompting or data augmentation may not be enough. Since we do not introduce new factual information in the hesitation text, the improvement from the hesitation-prompted input might be limited. Therefore, we resort to what difference hesitations can bring during decoding. Different from previous work Li et al. (2023c); Chuang et al. (2023); O’Brien and Lewis (2023) that contrasts between models, we use contrastive decoding (Li et al., 2023c) from the perspective of data:
$p_{\mathbf{CD}}(Y|X)=\text{softmax}(\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)})$ (5)
@@
With $p_{\mathbf{H}}(Y|X)$, the confidence difference could be noticed, especially when the prediction probability is low. While for the high-confidence tokens, the probability change brought by hesitations might be marginal. $p_{\mathbf{CD}}(Y|X)$ could be very uniform. Therefore, LLMs can mainly follow the base probability $p_{\theta}(Y|X^{\prime})$ to make predictions.
@@ 4 Experiments
We conducted experiments on five tracks of three hallucination evaluation benchmarks.
@@
Table 1: Truthfulness scores (%) on the three benchmarks. We also underline the second-best scores SH2 achieved. For ITI (Li et al., 2023b), "*" means we report the results on TruthfulQA from their paper since they use 5% of TruthfulQA questions in few-shot prompting. For 13b-CD, "*" means we report the results of Contrastive Decoding (CD) (Li et al., 2023c) from Chuang et al., 2023. They use LLaMA-7b as the amateur model and LLaMA-13b as the expert model. We maintain the same experimental settings for the other baselines and our SH2 in our implementation.
@@ 4.1 Benchmarks
TruthfulQA (Lin et al., 2022) is a benchmark to measure the truthfulness of a language model in question answering. It has 817 samples and both tracks of generation and discrimination for evaluation. To automatically evaluate the generation quality of LLMs, it introduces GPT-judge, a fine-tuned GPT-3. We use “Truth” to represent the percentage of the generated answers that are true and “Truth*Info” for answers that are both true and informative.
For the discrimination track, it offers sets of true and false reference answers for each question. We compute the likelihood of each answer given the question, and compare the probabilities of true answers against false answers to derive MC1, MC2 and MC3 scores. The definition of the MC (Multiple-Choice) metrics can be referred to in Appendix B.
@@
FACTOR (Muhlgay et al., 2023) puts more attention on the consistency of contexts and measures the tendency of a language model to generate factual information. It is a text completion task to identify the correct completion from the non-factual statements given the prefix. It contains two datasets of different sources: Wiki-FACTOR and News-Factor. There are 2994 and 1036 examples in each dataset. We gauge the factuality by whether the language model assigns the highest likelihood to the factually correct completion over the other options.

### v4

@@ Abstract
Large language models (LLMs) demonstrate great performance in text generation. However, LLMs are still suffering from hallucinations. In this work, we propose an inference-time method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 is based on a simple fact rooted in information theory that for an LLM, the tokens predicted with lower probabilities are prone to be more informative than others. Our analysis shows that these low-confidence tokens are more likely to be closely related to factual information, such as nouns, proper nouns, and adjectives. Therefore, we propose to “highlight” the factual information by selecting key tokens with the lowest probabilities and concatenating them to the original context, thus forcing the model to repeatedly read and hesitate on these tokens before generation. During decoding, we also adopt contrastive decoding to emphasize the difference in output probabilities brought by the hesitation. Experimental results demonstrate that our SH2, requiring no additional data or models, can effectively help LLMs elicit factual knowledge and distinguish hallucinated contexts by themselves. Significant and consistent improvements are achieved by SH2 for LLaMA-7b, LLaMA2-7b and Mistral-7b on various hallucination tasks. [^1] We release our code at https://github.com/LUMIA-Group/SH2.
@@
Jushi Kai  Tianhang Zhang  Hai Hu  Zhouhan Linthanks:  Corresponding author Shanghai Jiao Tong University json.kai@sjtu.edu.cn  lin.zhouhan@gmail.com
@@ 1 Introduction
Depending on massive training corpora, large language models (LLMs) have made tremendous progress in natural language understanding and text generation (Touvron et al., 2023a, Jiang et al., 2023, OpenAI, 2023). However, during reasoning and generation, LLMs could suffer from hallucinations and generate non-factual answers (Zhang et al., 2023).
To clear these falsehoods, some researchers construct datasets with higher quality and train LLMs to respond in the correct form (Taori et al., 2023, Zhou et al., 2023). But in the domain not included in the training, LLMs assign similar probabilities to correct and wrong choices since they do not have enough relevant knowledge to distinguish them.
To fill the gap of knowledge, retrieval augmentation methods Peng et al. (2023); Gao et al. (2023); Gou et al. (2023) leverage external knowledge bases and tools to correct the output of LLMs. Although they provide additional information for LLMs, it still seems impossible to guarantee zero error in the external knowledge. Other researchers propose decoding reformulation methods Li et al. (2023b); O’Brien and Lewis (2023); Chuang et al. (2023) to address hallucinations. However, most of them rely on external data with human labor or larger models to rescale probability distribution during decoding.
Recent works Wei et al. (2022); Goyal et al. (2023) have demonstrated that a few more computation steps for LLMs can make a difference. Wei et al. (2022) propose chain-of-thought (COT) prompting to elicit intermediate reasoning steps of LLMs. These intermediate steps can lead to better answers. However, elaborate prompts need to be prepared in advance. Goyal et al. (2023) append learnable pause tokens to the input prefix and train models from scratch. These delays introduced by pause tokens provide the model with more computation steps to generate better answers.
In this paper, we seek to leverage the LLM itself to highlight informative tokens and digest them as input during the hesitation steps to elicit truthful knowledge inside LLMs. We propose a simple yet effective method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 introduces hesitations to give LLMs more time to understand contexts and answer questions. For LLMs, the tokens assigned with lower probabilities are harder to predict, while more likely to be informative. LLMs can select these key tokens by themselves from the input and hesitate on these highlighted tokens. We calculate the difference brought by highlighted tokens through contrastive decoding (Li et al., 2023c) and integrate it into the output probability. Experiments on multiple tasks demonstrate that such a difference could elicit factual knowledge inside the model and successfully mitigate hallucinations. Unlike other methods, our method does not leverage any other external tools or data. Additionally, it can be directly deployed during inference with no more training.
@@ 2 Related Work
2.1 Hallucination Mitigation
Recent works to mitigate the hallucination of LLMs can be summed up into three categories.
Supervised Fine-Tuning
Many researchers pay attention to the curation of the training data and attempt to mitigate hallucinations through additional fine-tuning. Alpaca Taori et al. (2023) have collected 52K instruction-following data of massive tasks and fine-tuned the LLaMA-7b model Touvron et al. (2023a). Such an instruction tuning process is also known as supervised fine-tuning (SFT). Zhou et al. (2023) construct 1000 SFT samples with human labor for alignment and suggest that almost all knowledge in LLMs has been learned during pretraining. Moreover, Chen et al. (2023) leverage ChatGPT to automatically select high-quality data from Alpaca. These training approaches have high requirements for computational resources.
Retrieval Augmentation
@@ 2.1 Hallucination Mitigation
Retrieval augmentation approaches resort to external knowledge bases and tools to help correct hallucinations. Additional information is retrieved to provide relevant knowledge for LLMs and support their generation. Peng et al. (2023) and Gao et al. (2023) leverage search engines to attribute and refine the output of language models. Gou et al. (2023) enable multiple tools to correct responses of LLMs autonomously during the interaction with external tools.
Decoding Reformulation
@@
These approaches work on reformulating the probability distribution of outputs. Li et al. (2023b) introduce Inference-Time Intervention (ITI) to locate truthful directions of TruthfulQA Lin et al. (2022) and shift model activations toward truthfulness during inference. However, they need the data of TruthfulQA to train a domain-specific classifier for each attention head. Li et al. (2023c) propose Contrastive Decoding (CD) to capture the likelihood difference between large and small models. The difference signals which input texts should be preferred during decoding. O’Brien and Lewis (2023) utilize CD to improve reasoning quality for LLMs. Furthermore, DoLa Chuang et al. (2023) use the last layer as the expert model and the premature layer as the amateur, and contrast prediction probabilities between them. Different from these model-based CD strategies, our proposed method diverges by contrasting probabilities from a data-based perspective, offering a novel angle for decoding reformulation.
2.2 More Computations for Decoding
@@
The idea of using extra decoding steps when predicting hard tokens can be dated back to Adaptive Computation Time proposed by Graves (2017). Shin et al. (2020) introduce AUTOPROMPT to combine original inputs with trigger tokens to elicit knowledge from pretrained models. Wiegreffe et al. (2022) investigate that additional steps to generate rationales could lead to a more faithful model. More recently, Goyal et al. (2023) have also demonstrated that for LLMs, inserting extra computation steps to allow the model to pause before generation can enhance the performance on question answering and reasoning tasks. Nevertheless, their method only works when the model is both pre-trained and finetuned with pauses.
@@
Figure 1: The pipeline to construct and leverage our Self-Highlighted Hesitation. The original input X consists of the instruction, document and summary. The hesitation of key tokens is appended to the document in the hesitated input X’.
@@ 3 Self-Highlighted Hesitation
The illustration of our Self-Highlighted Hesitation is shown in Figure 1. For the original inference procedure, the instruction, document, and summary are directly fed into the LLM to generate the judgment. The LLM could be easily confused by the hallucinated context.
For our SH2, we underline the key tokens of the input document by the prediction likelihood. These tokens are listed as a hesitation and appended to the document. The prediction probability is scaled with the difference of confidence between the hesitated input X’ and the original input X. Our method can effectively help LLMs identify the hallucinated context. We will elaborate our method in the following subsections.
@@ 3.2 Relation between Key Tokens and Factual Information
where $POS(\cdot)$ is the function to derive POS tags and #$\#\{POS(X_{i})=z_{k}\}$ is to count how many words in $X_{i}$ have the POS tag of $z_{k}$.
For a dataset with $m$ documents, the normalized recall of POS tag $z_{k}$ can be calculated by:
@@
where $T(X_{i},\eta)$ is the set of words that are among the lowest $\eta$ portion of probability predicted by the language model in $X_{i}$. The numerator measures the frequency of the POS tag $z_{k}$ in the subset of the lowest probability. The $\eta$ in the denominator is to normalize the scale of $\Delta_{\eta}(z_{k})$ with different $\eta$.
$\Delta_{\eta}(z_{k})$ is the frequency difference between subsets and documents. It measures how hard words with the POS tag $z_{k}$ are for LLMs to predict. Figure 2 demonstrates the normalized top-$\eta$ recall of the most frequent POS tags for the summarization track of HaluEval Li et al. (2023a).
Larger $\Delta_{\eta}(z_{k})$ means that $z_{k}$ is more concentrated in the hardest part of documents. For example, although there are 10 times as many prepositions (IN) as superlative adjectives (JJS) in the set of words with the lowest 1% ($\eta=1$) portion of probability, the recall of IN is even smaller. This is because IN is more frequent naturally. There are 60 times as many IN as JJS throughout documents, but IN is less concentrated in the hardest part.
It can be observed from Figure 2 that content words such as adjectives (JJ), nouns (NN), proper nouns (NNP), adverbs (RB) and conjugated verbs (VBD, VBG, VBP) are more difficult to predict and more concentrated in the hardest part, as the light color in the heatmap indicates. These content words usually contain factual information. On the other hand, function words such as conjunctions (CC), determiners (DT), and prepositions (IN) are less informative and less concentrated in the hardest part. [^2] The base form of the verbs (VB) seems to be an exception. It is probably because VBs have less information load than conjugated verb forms, as the latter also encode tense, aspect, and other information about the verb. Therefore, picking out key tokens that carry more content, rather than those that serve as grammatical functions, could potentially help the model focus on factual information.
@@ 3.3 Construction of Hesitations
To this end, we attempt to solve hallucinations by highlighting key tokens that are hard to predict. For the text sequence $X=(x_{0},x_{1},\dots,x_{n})$, we can obtain the corresponding probability sequence by the language model:
@@
We use the following strategy to select key tokens according to their probability $\hat{P}(X)$ and construct hesitations with these tokens:
First, construct a candidate key-token set by selecting $\eta\cdot n$ tokens with the lowest prediction probabilities of $\hat{P}(X)$, where $\eta\in(0,1)$ is the sampling proportion we preset for $X$.
For long documents, to avoid the key-token set getting dominated by one or two POS tags, we introduce the drop-out rate $\lambda\in[0,1]$ to randomly select key tokens among a larger pool of candidates, whose size is determined by $\eta$. As a result, $(1-\lambda)\cdot\eta\cdot n$ tokens are retained in the target key-token set $T(X)$ to fill as input during hesitation steps.
Finally, construct the Hesitation $H(X)$ with the target key-token set for the input context: $H(X)=``Pondering:<T(X)>."$, where ``$Pondering:"$ is the prefix. We keep the tokens’ order of $T(X)$ in which they appear in the original text.
@@ 3.4 Contrastive Decoding on Hesitations
Using hesitations as prompting or data augmentation may not be enough. Since we do not introduce extra factual information, the improvement from the hesitated input might be limited. Therefore, we resort to what difference hesitations can bring during decoding. Different from previous work Li et al. (2023c); Chuang et al. (2023); O’Brien and Lewis (2023) that contrasts between models, we use contrastive decoding (Li et al., 2023c) from the perspective of data:
$p_{\mathbf{CD}}(Y|X)=\text{softmax}\left(\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)}\right)$ (5)
@@
With the base probability scaled by the contrastive term, the confidence difference could be noticed. For tokens with low prediction probabilities, $p_{\mathbf{H}}(Y|X)$ is dominated by the contrastive term. For high-confidence tokens, the probability change brought by hesitations might be marginal. $p_{\mathbf{CD}}(Y|X)$ could be very uniform. Therefore, LLMs can mainly follow the base probability $p_{\theta}(Y|X^{\prime})$ to make predictions. An illustration is given in Appendix D.1.
@@ 4 Experiments
We conduct experiments on five tracks of three hallucination evaluation benchmarks.
@@
Table 1: Truthfulness scores (%) on the three benchmarks. The second-best scores for the LLaMA-7b backbone are also underlined. For ITI, "*" means we report results on TruthfulQA from their paper since they trained a probe for inference. For 13b-CD, "*" means we report results of Contrastive Decoding from Chuang et al., 2023. They use LLaMA-13b as the expert model. We maintain the same experimental settings for SH2 and other baselines of our implementation. Mistral-7b is not evaluated on the generation track of TruthfulQA due to API problems of OpenAI.
@@ 4.1 Benchmarks
TruthfulQA (Lin et al., 2022) is a benchmark to measure the truthfulness of a language model in question answering. It has 817 samples for generation and discrimination tracks. To automatically evaluate the generation quality of LLMs, it introduces GPT-judge, a fine-tuned GPT-3. We use “Truth” to represent the percentage of truthful answers and “Truth*Info” for generated answers that are both true and informative.
For the discrimination track, it offers sets of true and false reference answers for each question. We compute the likelihood of each answer given the question, and compare probabilities of true answers against false answers to derive MC1, MC2 and MC3 scores. The definition of the MC (Multiple-Choice) metrics can be referred to in Appendix B.
@@
FACTOR (Muhlgay et al., 2023) puts more attention on the consistency of contexts and measures the tendency of language models to generate factual information. It is a text completion task to identify the correct completion from non-factual statements given the prefix. It contains two datasets of different sources: Wiki-FACTOR and News-Factor. There are 2994 and 1036 examples in each dataset. We gauge the factuality by whether the model assigns the highest likelihood to the factually correct completion over the other options.